Say that you purchase a house for $248,000 by getting a mortgage for $220,000 and paying a $28,000 down payment. If you get a 30-year mortgage with an interest rate of 8 percent, what are the monthly payments? (Do not round intermediate calculations and round your final answer to 2 decimal places.)

Find Payment

What would the loan balance be in ten years? (Round the payment amount to the nearest cent but do not round any other interim calculations. Round your final answer to 2 decimal places.)

Find Loan Balance

If the house appreciates at 4 percent per year, what will be the value of the house in ten years? (Do not round intermediate calculations and round your final answer to 2 decimal places.)

Find Future Value

How much of this value is your equity? (Do not round intermediate calculations and round your final answer to 2 decimal places.)

Find Equity
Answer #1

Calculating Monthly Payment on Loan,

Using TVM Calculation,

PMT = [PV = 220,000, FV = 0, N = 360, I = 0.08/12]

PMT = $1,614.28

Calculating Loan Balance in 10 years,

Using TVM Calculation,

FV = [PV = 220,000, PMT = -1,614.28, N = 120, I = 0.08/12]

FV = $192,994.73

Value of House in 10 years = 248,000(1.04)10 = $367,100

Equity in 10 years = 367,100 - 192,995

Equity in 10 years = $174,105